OpenStudy (poison_princess1):

Silver has a density of 10.5g/cm3 and gold has a density of 19.3g/cm3.Which would have the greater mass, 5 cm3 of silver or 5cm3 of gold?

OpenStudy (poison_princess1):

I NEED HELP

OpenStudy (mhchen):

Sooooooooooooooooooooooooo Density = mass / volume Silver: \[10.5~g/cm^3 = \frac{mass}{5~cm^3}\] Gold: \[19.3~g/cm^3 = \frac{mass}{5~cm^3}\] Now solve for mass on both equation, you know how to do that right?

OpenStudy (poison_princess1):

Not really

OpenStudy (will.h):

By multiplication

OpenStudy (poison_princess1):

With the density?

OpenStudy (will.h):

Mass = Density * volume So do the calculations and find the results

OpenStudy (poison_princess1):

Ok

OpenStudy (poison_princess1):

It will be for silver will be 52.5 and for gold it will 96.5 right ?

OpenStudy (poison_princess1):

Then I divided that with the volume right

OpenStudy (poison_princess1):

So really the question is ask the mass right?